OpenStudy (anonymous):

a closed flask contains a 0.25 moles of O2 which exerts a pressure of 0.50 atm. if 0.75 moles of CO2 is added to the container what is the total pressure in the flask

OpenStudy (ivancsc1996):

Well you need ideal gas law. So the equation is:\[\frac{ P _{1}V _{1} }{ n _{1}T _{1} }=\frac{ P _{2}V _{2} }{ n _{2}T _{2} } \rightarrow P _{2}=\frac{ P _{1}V _{1}n _{2}T _{2} }{ n _{1}T _{1}V _{2} }\]So if you cancel the Ts and the Vs since the temperature and the volume is equal you get that \[P _{2}=P _{1}\frac{ n _{2} }{ n _{1}} =0.5atm \frac{ 0.75mol }{ 0.25mol }=1.5atm\]